OpenStudy (anonymous):

Give on example of an equation with variables on both sides that is an identity and one equation with variables on both sides that has no solution. Justify your examples by solving the equations.

OpenStudy (anonymous):

identity: x + 5 = x + 5 this becomes x-x = 5-5 0 = 0 identity.. no solution : x+6 =x+5 x-x = 5-6 0 = -1 no solution.